OpenStudy (anonymous):

1. A car is moving on a straight road with a uniform acceleration. The following table gives the speed of the car at various instants of time. Time (s) 0 10 20 30 40 50 Speed (ms-1) 5 10 15 20 25 30 (i) Draw the shape of speed-time graph representing the above sets of observations.(1) (ii) Find the acceleration of the car. (1) (iii) Distance travelled between 30sec to 50 sec. [2]

OpenStudy (anonymous):

uniform acc means no change in v wth rspct to time so take velcty on x axs nd time on y axs nd draw a graph. then chose any two values of v nd t nd find acc. ans of 3rd que can b filled by formula \[ vf -vi \div t2-t1\]

OpenStudy (ganpat):

 accelaration = (v-u)/ t distance travelled = x = ut + 1/2 * at2..